OpenStudy (anonymous):

Solve the system algebraically. 4x - 5y = 3 x = 3 - 1/2y What is the value of x? -------------- 9/7 24/14 33/14

OpenStudy (anonymous):

solve for y by substitution and evaluate --> 4(3 - 1/2y) - 5y = 3

OpenStudy (anonymous):

then substitute the y value back into one of the equations to solve for x

OpenStudy (anonymous):

4x-5y=3...A x = 3 - 1/2y or x+1/2y=3...B mutiply the second equation with -10 -10(x+1/2y=3) -10x-5y=-30 then use both equation A and B eliminate the common 4x-5y=3 -10x-5y=-30 + + = + 14x=33 (5y will be cancel out) x=33/14 x=2.35

OpenStudy (anonymous):

So it C is the right answer i.e 33/14

OpenStudy (anonymous):

Awesome thanks!

OpenStudy (anonymous):

u r welcum =)